OpenStudy (anonymous):

The expression 3^1/3 divided by 3^ -2/3 is equivalent to ?? is the answer 3?

OpenStudy (shamil98):

\[\huge \frac{ 3^{1/3} }{ 3^{-2/3} }\] Yes, you are correct. When you divide exponents you subtract them: \[\huge \frac{ x^m }{ x^n } = x^{m-n}\]

OpenStudy (kc_kennylau):

yes

OpenStudy (anonymous):

Thank youuuu!!! :D

OpenStudy (shamil98):

\[\huge 3^{1/3 -(-2/3)} = 3^{3/3} = 3^1 = 3\]